OpenStudy (anonymous):

Can someone take me through the steps of the proof of Tan(theta/2)= plus/minus sqrt((1-cos(theta)/(1+cos(theta))= (1-cos(theta))/sin(theta)?

OpenStudy (anonymous):

Basically take me through \[\tan \frac{\theta}{ 2 }=\pm \sqrt{\frac{ 1-\cos{\theta} }{ 1+\cos{\theta} }}=\frac{ 1-\cos{\theta} }{ \sin{\theta}}\]

OpenStudy (anonymous):

@primeralph

OpenStudy (anonymous):

@goformit100 @mathslover @xlegendx

OpenStudy (anonymous):

hmm half-angle formulas huh. do you know the half angle formula for cos and sin?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Oh, i feel dumb, is it becuase since they're both square roots that you can put them under the same radical and the 2's in the denominator cancel out?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

oh, alrighty, then